# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v4

@@ 5.2.1 Explainability in Multi-Agent Decision Making
Explainability refers to making the inner workings and decisions of AI systems interpretable to humans [145]. In AMAS settings, outcomes emerge from the interactions among specialized agents rather than a single model prediction [146]. Therefore, explainability must cover both individual agent decisions and the inter-agent dynamics that lead to emergent results. Recent research emphasises that user trust hinges on understanding how each agent contributes and how their collaboration is orchestrated [111]. Traditional single-model XAI techniques often ignore this multi-agent dimension, so new methods are required.
Challenges Explaining multi-agent decisions requires tracing which agent performed which sub-task, what information they used, and how their interactions influenced the final outcome [56]. Additional challenges include: (1) Emergent behaviour: interactions among agents can lead to behaviour that was not explicitly programmed; explaining these emergent patterns requires capturing interaction chains.; (2) Verification of intermediate steps: vanilla CoT prompting in LLMs often produces plausible-sounding but unchecked reasoning. Without verification, small errors can propagate
Recent directions Recent works introduce new frameworks and techniques for multi-agentic explainability:
  • Layered Chain-of-Thought prompting: Instead of producing a single monolithic reasoning trace, Layered-CoT divides reasoning into discrete layers [111]. Each layer answers a sub-question, and a specialized Reasoning Agent generates the partial chain-of-thought while Verification Agents and User-Interaction Agents cross-check the intermediate steps against knowledge graphs or human feedback. This agent-based decomposition localizes errors, prevents contradictions, and ensures that each partial solution is validated before the model proceeds.
  • Integrated multi-agent workflows: Multi-agent workflows embed explainability into the process itself rather than as a post-hoc add-on. For example, in a DSM-5 mental-health diagnosis workflow, one agent collects patient information while a diagnostician agent retrieves relevant DSM-5 criteria and explicitly links conversational utterances to diagnostic rules. This integration produces transparent, step-by-step diagnostic predictions and shows why specific recommendations are made [147]. The study notes that step-by-step reasoning improves perceptions of fairness and reliability and that embedding a diagnostician agent moves beyond general synthetic dialogues by providing a self-documenting rationale.
  • Modular resume-screening systems: In recruitment, a multi-agent architecture divides the process into extraction, evaluation, summarization and formatting. The evaluation agent dynamically retrieves company-specific hiring criteria via Retrieval-Augmented Generation (RAG), and each stage remains clearly defined. This modularity allows recruiters to trace how a candidate was assessed and why a particular score was assigned [148].
  • Explainable credit assignment in multi-agent reinforcement learning: A centralized LLM critic can provide individualized credit assignments to agents and explain its feedback strategy. Unlike earlier deep RL methods, this approach uses the language-based critic to justify why each agent receives a particular reward and to show how it decomposes the global reward [149].
Approaches for explainability in multi-agent LLM systems The following techniques adapt traditional XAI methods to the multi-agent setting:
@@
  • Feature Attribution (LIME/ SHAP) for agents: Techniques like LIME [112] or SHAP [113] can be extended to attribute an agent’s output to specific input components (e.g., particular context tokens or inter-agent messages). Highlighting which parts of the input or which inter-agent signals most influenced an agent’s action helps illuminate why a specific agent behaved as it did [150].
  • Counterfactual and Causal Analysis Beyond local feature attributions, “what-if” analyses are increasingly important for multi-agent explainability. Counterfactual techniques examine how the system’s behavior would change if certain conditions or agent behaviors were altered [151]. For example, one can remove or modify a particular agent’s contribution and then observe how the collective outcome changes. This approach, rooted in causal inference, helps isolate each agent’s role in collaborative decision-making.
  • Reasoning Trace Logging: Instead of treating the agents as black boxes, an explicit trace of their intermediate reasoning can greatly enhance transparency. Logging the step-by-step reasoning process – for example, an agent CoT prompts or the dialogue exchanged between agents – creates an audit trail of how the decision emerged [75]. By examining this trace, one can see how agents influence each other, which agent contributed what, and how partial conclusions build up to the final decision. This visibility into the multi-agent process addresses a key blind spot of standard XAI, effectively “opening up” the black box of inter-agent interactions.
@@
  • Natural Language Explanations: Multi-agent systems can include a separate explainer agent that synthesizes logs and feature attributions into human-readable narratives. This agent (often another LLM) translates the technical reasoning traces into conversational explanations that answer questions like “Why did the agents choose action X?” Integrated workflows, such as DSM5AgentFlow [147], demonstrate that embedding such explanation directly in the pipeline improves transparency and trust.
  • Role-oriented modular architectures: Assigning clear roles to agents (e.g., reasoning, verification, summarization, credit assignment) enhances explainability because each stage of decision-making is distinct and can be inspected independently [148]. Modular designs also allow domain experts to adjust or replace specific agents without affecting the entire system.
  • Explainable reward critics: In multi-agent reinforcement learning, LLM-based critics can decompose global rewards and provide natural-language justifications, making credit assignment transparent [149] . This approach helps developers understand how credit is distributed across agents and why certain behaviors are encouraged or discouraged.
@@ 5.2.2 ModelOps : Lifecycle Management for LLM Agents
ModelOps (Machine Learning Operations) extends the MLOps paradigm beyond model training and deployment to include governance, risk management and continuous oversight for all types of AI models [131]. Unlike MLOps, ModelOps manages not only individual models but also composite systems of agents and heterogeneous models across different business units.
ModelOps provides end-to-end oversight for AI in production by enforcing model governance and compliance: tracking lineage, ownership, and metadata to meet regulations such as GDPR and HIPAA , and by using policy-driven deployment gates so only approved models go live, with controlled roll-outs and roll-backs [132]. It continuously monitors risk and bias, detecting drift, fairness issues, hallucinations, and robustness regressions. ModelOps also manages heterogeneous assets: rule-based systems, LLMs, and composite multi-agent architectures; ensuring consistent visibility and lifecycle control across the portfolio. Finally, it integrates tightly with enterprise IT, wiring AI workflows into CI/CD, IT service management (ITSM), identity and access management (IAM), and cloud-security tooling [132]. For generative AI and AMAS, ModelOps must handle prompt versioning, guardrails and content filtering enforcement, fine-tuning lineage tracking, usage and cost observability, and ethical compliance. Effective ModelOps thus requires:
  • Version control and CI/CD pipelines: Track versions of each agent’s model and prompt configuration. Automated pipelines should test safety, performance and cost impacts whenever an agent’s logic or model is updated, using multi-agent simulations to validate that new behaviours do not introduce regressions or unsafe interactions.
  • Hierarchical monitoring in multi-agent frameworks: Systems such as MegaAgent [128] decompose tasks into subtasks and dynamically form groups of agents. They employ hierarchical monitoring, where each agent tracks its actions with a checklist, group-level administrators oversee their agents, and system-level supervisors review outputs. This layered monitoring ensures that changes to one agent do not compromise the entire system.
  • Risk measurement and oversight policies: The NIST generative AI profile recommends documenting risks and updating measurement approaches regularly; it suggests using gradient-based attributions, occlusion and counterfactual prompts to improve explainability and transparency, and establishing policies for oversight across the lifecycle [126].
@@ 5.2.3 Application Security for LLM Agents
Security in AMAS systems must address vulnerabilities such as prompt injection, data exfiltration, and the misuse of external tools. According to OWASP [137, 138], prompt injection manipulates an LLM through malicious instructions, potentially causing the model to leak data, execute unauthorized commands or make harmful decisions. Multimodal models face cross-modal injection where malicious content in one modality influences the output of another. To mitigate these risks:
  • Prompt hygiene and hardening: Sanitize and filter user inputs, apply secure prefixes, and validate instructions before they reach an agent. Constraining the output format and employing guardrails can limit the model’s response surface [38].
  • Defense-in-depth design patterns: Existing techniques such as adversarial training and user confirmation provide partial protection, but research shows they do not guarantee safety. Design patterns [133] like Action-Selector and Plan-Then-Execute restrict an agent’s capabilities after processing untrusted input; the agent selects from a predefined list of allowed actions or produces a plan that cannot be altered by tool outputs, thereby isolating untrusted data from execution paths [58].
  • Least-privilege access and authentication: Enforce strong authentication for both human users and agents and apply role-based access controls so agents can only invoke the minimum set of tools necessary to perform their task. Continuous monitoring should detect anomalous requests; if an agent deviates from normal patterns, automated systems should flag or suspend it [152].
  • Post-processing and anomaly detection: Implement filters to detect and remove sensitive or inaccurate content from outputs and regularly retrain models to reduce hallucinations. Maintain robust logging and retention policies so that investigators can audit interactions without exposing private data.
Recent frameworks also emphasis hierarchical monitoring and cross-checking among agents to detect inconsistencies [128]. LangChain [153], AutoGen [98], and CrewAI [87] introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise. Training agents with adversarial examples [134] and red-teaming [154] scenarios further strengthens robustness against malicious inputs.
@@ 5.2.4 Model Privacy and Data Protection
AMAS also poses unique privacy challenges because agents often share information via memory stores and message passing. Without strict controls, an agent may inadvertently leak sensitive data [40]. A layered privacy strategy should include:
  • Differential Privacy (DP): Adding calibrated noise to training data prevents individual data points from being re-identified and reduces the risk that models memorize or regurgitate personal information [139]. DP should be assessed on a case-by-case basis, since heavy noise can harm accuracy.
  • Data anonymization and minimization: Apply robust anonymization and pseudonymization techniques and regularly test them for effectiveness [140]. Limit data collection to what is strictly necessary for the agent’s task and enforce strict retention policies.
  • Secure Multi-Party Computation (MPC): MPC allows parties to jointly compute a function without revealing their private inputs. In privacy-preserving LLM services, secure MPC can protect user prompts from the model provider; for example, the PUMA framework uses MPC to ensure that prompts remain hidden during inference, while maintaining reasonable performance (it can evaluate a 7-billion-parameter model within minutes) [141].
  • Homomorphic Encryption (HE): HE enables computation on encrypted data and offers cryptographically provable privacy. Recent work demonstrates an encryption-friendly LLM architecture that uses LoRA fine-tuning and Gaussian kernels to speed up fully homomorphic encryption (FHE) operations, allowing encrypted inference with performance comparable to plaintext models [142].
  • Trusted Execution Environments (TEEs): TEEs provide hardware-enforced isolation that protects code and data [143]. They partition processors into secure and normal worlds; implementations include Intel SGX, AMD SEV and Arm TrustZone. In secure LLM inference, TEEs ensure that models run only inside authenticated enclaves, with attestation verifying model integrity.
@@
These techniques, combined with careful consent management and compliance with local regulations, help maintain confidentiality when agents share or process sensitive information.
@@
Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art. 13–14) [125] and the NIST AI RMF trustworthiness characteristic “Explainability & Interpretability” [3]; also supported by NISTIR 8312 (Four Principles of Explainable AI) [155]. Promotes well-justified (rationale) and human-interpretable outputs with traceable logs [119]. ModelOps: Anchored in the NIST AI RMF “Govern/Manage” functions [3] and ISO/IEC 42001:2023 (AI management system) [156]; secure SDLC controls via NIST SP 800-218/218A [^5] NIST, 2024. Emphasizes lifecycle gates, model versioning, rollback policies, and CI/CD with safety gates. Risk Management & Impact Assessment: EU AI Act (Art. 9, 72–73) [125], ISO/IEC 23894 (AI risk management) [^6] ISO/IEC 23894, and ISO/IEC 42005:2025 (AI impact assessment) [129]. Iterative hazard analysis, impact registers, incident response, and disclosure playbooks. Data Governance & Quality: EU AI Act (Art. 10) [125] mapped to the ISO/IEC 5259 series (data-quality model, measures, processes, governance) [^7] ISO/IEC 5259 series. Data minimization, lineage, coverage/bias checks, and sampling documentation. Robustness & Safety: EU AI Act (Art. 15) [125] and ISO/IEC 24029-1 (robustness assessment) [32]. Adversarial/perturbation testing, fault tolerance, fallback/kill-switch triggers, and rollback criteria. Application Security: Guided by OWASP LLM Top-10 [137] and OECD AI Principles on robustness [127]. Access controls, prompt/input mediation, sandboxing, egress controls, and auditable trails. Model Privacy: Aligned with GDPR (Art. 25) [34], HIPAA §164 [35], and OECD AI Principles [127]. Encryption in transit/at rest, DP budgeting, data minimization, and access-audit pass thresholds. Documentation & Traceability: EU AI Act (Arts. 11–12; Annex IV) [125]. Complete technical documentation, event logs, model/data cards, decision records, and regulator-ready evidence packages. Post-Market Monitoring: EU AI Act (Arts. 72–73) [125]. Telemetry for drift/misuse, continuous evaluations, feedback loops, controlled recalls, and stakeholder notifications. Governance: Grounded in GDPR, CCPA [144], and OECD AI Principles [127]. Institutional oversight, policy compliance, role-based accountability, regulatory traceability, and human-in-the-loop for high-risk/multi-agent deployments. Fairness & Human Rights: OECD AI Principles [127] and ISO/IEC TR 24027 (bias) [^8] ISO/IEC TR 24027. Lifecycle bias controls, affected-party impact reviews, and mitigation sign-offs.
5.2.5 Governance
Governance integrates ethical, legal and societal considerations into the design and operation of AI systems. Regulations such as the EU AI Act [125], NIST AI RMF [3] and OECD AI [127] provide guidance that aligns with TRiSM pillars of Explainability, ModelOps, Security and Privacy.
  • EU AI Act - Human Oversight (Article 14): High-risk AI systems must be designed so that natural persons can effectively oversee them and prevent or minimize risks to health, safety or fundamental rights [1]. The AI system should enable overseers to understand capabilities and limitations, detect anomalies, interpret outputs, decide not to use the system, or stop its operation through a ‘stop’ button. For certain identification tasks, decisions must be independently verified by at least two competent individuals.
  • OECD AI Principles: These intergovernmental principles, updated in 2024, promote human-centric, trustworthy AI [127]. They emphasise transparency and explainability: AI actors should provide meaningful information about data sources, factors and logic to allow affected individuals to understand and challenge outputs. Robustness, security and safety require AI systems to function appropriately under normal and adverse conditions and include mechanisms to override or decommission systems to prevent harm. Accountability means AI actors are responsible for the proper functioning of AI systems; they must ensure traceability of datasets and decisions and apply systematic risk management throughout the lifecycle.
  • NIST AI Risk Management Framework (AI RMF): The AI RMF and its Generative AI profile outline four core functions: govern, map, measure and manage [3]. For generative AI, NIST recommends documenting risks and potential impacts, applying interpretability techniques (gradient-based attributions, occlusion, counterfactual prompts) to improve transparency, and establishing policies for oversight across the AI lifecycle.
  • Auditability and role-based access: Agentic AI systems should maintain auditable logs of agent actions, decisions and data flows and assign clear roles (e.g., reasoning, verification, summarisation) so that responsibilities are transparent [157]. Role-based access controls help enforce accountability and prevent unauthorized agents from accessing sensitive functions.
  • Global harmonisation: Many jurisdictions (e.g., GDPR, CCPA) require data subject rights, consent and privacy by design. Frameworks like the OECD Principles and NIST AI RMF promote international cooperation and interoperability.
These governance frameworks (as illustrated in Figure 6) underscore the need for transparent accountability structures, human oversight, robust privacy controls, and documented processes in AMAS. TRiSM provides a practical scaffolding for operationalizing these requirements, aligning lifecycle management, security, privacy and governance to build trustworthy Agentic AI.
@@ 5.2.5 Governance
6 Evaluation of Agentic AI Systems
Agentic AI systems demand evaluation beyond traditional accuracy. We outline five categories: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct facet of performance and real-world impact. Compared to a related work [158], which focuses on proportional impacts across general AI, our work targets Agentic AI and links risks to their causal sources (e.g., prompt injection, memory poisoning). We further introduce an evaluation template with new metrics the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE) to capture multi-agent dynamics. A summary appears in Table 5.

### v5

@@ 4.2.1 Explainability in Multi-Agent Decision Making
In AMAS, explainability is less about interpreting a single model and more about understanding how decisions emerge from interacting agents [yossef2024explainable]. In AMAS settings, outcomes often emerge from interactions among specialized agents rather than from a single model prediction [rosenfeld2019explainability]. As a result, explainability must account for both (i) individual agent decisions and (ii) the inter-agent dynamics that produce emergent system-level behavior. Recent work emphasizes that user trust depends on understanding each agent’s contribution and how collaboration is orchestrated [sanwal2025layered]. Traditional single-model XAI methods often overlook this multi-agent dimension, motivating new approaches tailored to AMAS.
Challenges. Explaining multi-agent decisions requires tracing which agent performed which sub-task, what information it used, and how inter-agent exchanges influenced the final outcome [de2025open]. Key challenges include:
  • Emergent behavior: agent interactions can produce behaviors that were not explicitly specified; explaining these patterns requires capturing and summarizing interaction chains.
  • Verification of intermediate steps: vanilla CoT prompting can yield plausible but unverified reasoning. Without systematic verification, small errors may propagate across agents and compound into incorrect final decisions.
Approaches for explainability in AMAS The following techniques adapt classical XAI ideas to AMAS:
@@
  • Feature attribution (LIME/SHAP) for agents: LIME [ribeiro2016should] and SHAP [lundberg2017unified] can be extended to attribute an agent’s output to specific inputs (e.g., context spans, retrieved evidence, or inter-agent messages), clarifying which signals most influenced an agent’s action [bilal2025llms].
  • Counterfactual and causal analysis: “What-if” analysis evaluates how outcomes change when an agent’s contribution or a key condition is modified [gyevnar2023causal]. This helps isolate agent roles and identify causal dependencies in collaborative decision-making.
  • Reasoning-trace logging: Logging intermediate reasoning artifacts (e.g., CoT traces, tool calls, and inter-agent dialogues) creates an auditable trail of how decisions emerged [yao2023reactsynergizingreasoningacting]. Such traces reveal which agent contributed what and how partial conclusions accumulate into final outputs.
@@
  • Natural-language explanations: An explainer agent can synthesize traces and attributions into human-readable narratives that answer questions such as “Why did the system choose action X?” Embedded explainers in workflows (e.g., DSM5AgentFlow) can improve transparency and trust [ozgun2025trustworthy].
  • Role-oriented modular architectures: Assigning explicit roles (e.g., reasoning, verification, summarization, credit assignment) improves inspectability because each stage is separable and reviewable [lo2025ai].
  • Explainable reward critics: In multi-agent RL, LLM-based critics can provide transparent reward decomposition and justification, supporting debugging and governance of agent incentives [nagpal2025leveraging].
Recent directions. Recent work proposes frameworks and techniques that embed explainability directly into multi-agent operation:
  • Layered Chain-of-Thought prompting: Layered-CoT decomposes reasoning into discrete layers, where each layer answers a sub-question [sanwal2025layered]. A specialized Reasoning Agent produces partial reasoning, while Verification and User-Interaction Agents cross-check intermediate outputs (e.g., against knowledge graphs or human feedback). This decomposition can localize errors, reduce contradictions, and promote validation before proceeding.
  • Integrated multi-agent workflows: Some systems integrate explanation into the workflow rather than adding it post hoc. For example, in a DSM-5 mental-health diagnosis workflow, one agent collects patient information while a diagnostician agent retrieves DSM-5 criteria and explicitly links conversational evidence to diagnostic rules, producing a self-documenting rationale [ozgun2025trustworthy].
  • Modular resume-screening systems: In recruitment, multi-agent pipelines separate extraction, evaluation, summarization, and formatting. The evaluation agent retrieves company-specific criteria via RAG, enabling recruiters to trace how candidate evidence maps to scores [lo2025ai].
  • Explainable credit assignment in multi-agent reinforcement learning: A centralized LLM critic can provide per-agent credit assignment and natural-language justification, explaining how global rewards are decomposed across agents [nagpal2025leveraging].
@@ 4.2.2 ModelOps : Lifecycle Management for LLM Agents
ModelOps (Machine Learning Operations) extends the MLOps paradigm beyond model training and deployment to include governance, risk management, and continuous oversight across AI systems [lefevre2022modelops]. In production, ModelOps provides end-to-end oversight by enforcing governance and compliance controls, including lineage, ownership, and metadata tracking to support regulatory requirements (e.g., GDPR and HIPAA), as well as policy-driven deployment gates so that only approved models go live with controlled rollouts and rollbacks [Sinha2025ModelOps]. It also enables continuous risk monitoring to detect drift, bias and fairness issues, hallucinations, and robustness regressions.
For AMAS, ModelOps must additionally support prompt and toolchain versioning, enforcement of guardrails and content filtering, fine-tuning lineage tracking, usage and cost observability, and ethical compliance. Effective ModelOps for LLM agents therefore requires:
  • Version control and CI/CD pipelines: Track versions of each agent’s model, prompts, tools, and configuration. Automated pipelines should test safety, performance, and cost impacts whenever agent logic changes, using multi-agent simulations to validate that updates do not introduce regressions or unsafe interactions.
  • Hierarchical monitoring in multi-agent frameworks: Systems such as MegaAgent [wang2025megaagent] decompose tasks into subtasks and dynamically form agent groups. They apply hierarchical monitoring, where agents log actions (e.g., via checklists), group-level administrators oversee agent teams, and system-level supervisors review final outputs. This layered oversight helps ensure that changes in one agent do not compromise system-wide behavior.
  • Risk measurement and oversight policies: The NIST generative AI profile recommends documenting risks, updating measurement approaches regularly, and establishing lifecycle oversight policies. It also highlights techniques such as gradient-based attributions, occlusion tests, and counterfactual prompting to strengthen transparency and explainability [ai2024artificial].
@@ 4.2.3 Application Security for LLM Agents
Security in AMAS must address vulnerabilities such as prompt injection, data exfiltration, and the misuse of external tools. OWASP identifies prompt injection as a core risk in LLM applications: attackers manipulate the model with malicious instructions, potentially causing data leakage, unauthorized tool use, or unsafe decisions [owasp2024llmtop10, owasp2025agentic]. Multimodal models further face cross-modal injection, where malicious content in one modality (e.g., text in an image) influences behavior in another. To mitigate these risks:
  • Prompt hygiene and hardening: Sanitize and filter user inputs, apply secure system prefixes, and validate instructions before they reach an agent. Constraining output formats and enforcing guardrails can reduce the attack surface [raza2025justhumansneedvaccines].
  • Defense-in-depth design patterns: Techniques such as adversarial training and user confirmation provide partial protection, but do not guarantee safety. Design patterns [beurer2025design] such as Action-Selector and Plan-Then-Execute restrict capabilities after untrusted input is processed: the agent selects from a predefined set of allowed actions, or produces a plan that cannot be altered by tool outputs, thereby isolating untrusted data from execution paths [qiao2024agent].
  • Least-privilege access and authentication: Enforce strong authentication for both users and agents, and apply role-based access control so agents can invoke only the minimum set of tools required for their task. Continuous monitoring should detect anomalous requests; if an agent deviates from expected behavior, automated systems should flag or suspend it [barbera2025ai].
  • Post-processing and anomaly detection: Apply output filters to detect sensitive or inaccurate content, and update or retrain models to reduce hallucinations. Maintain robust logging and retention policies to support audits and incident response while minimizing exposure of private data.
Recent frameworks also emphasize hierarchical monitoring and cross-checking among agents to detect inconsistencies [wang2025megaagent]. Tooling ecosystems such as LangChain [langchain2024function], AutoGen [autogen_agents_tutorial], and CrewAI [crewAI] operationalize verification through agent-level checks (e.g., reviewers/critics) and, in some cases, trust or reputation signals that encourage agents to validate each other’s outputs and flag potential compromise. Finally, adversarial training with attack examples [standen2025adversarial] and systematic red-teaming [ganguli_red_2022] can further strengthen robustness against malicious inputs.
@@ 4.2.4 Model Privacy and Data Protection
Without strict controls, an agent may inadvertently leak sensitive data [ko2025sevensecuritychallengessolved]. A layered privacy strategy should include:
  • Differential Privacy (DP): Adding calibrated noise to training data prevents individual data points from being re‑identified and reduces the risk that models memorize or regurgitate personal information [dwork2006calibrating]. DP should be assessed on a case‑by‑case basis, since heavy noise can harm accuracy.
  • Data anonymization and minimization: Apply robust anonymization and pseudonymization techniques and regularly test them for effectiveness [sweeney2002k]. Limit data collection to what is strictly necessary for the agent’s task and enforce strict retention policies.
  • Secure Multi‑Party Computation (MPC): MPC allows parties to jointly compute a function without revealing their private inputs. In privacy‑preserving LLM services, secure MPC can protect user prompts from the model provider; for example, the PUMA framework uses MPC to ensure that prompts remain hidden during inference, while maintaining reasonable performance (it can evaluate a 7‑billion‑parameter model within minutes) [lindell2009secure].
  • Homomorphic Encryption (HE): HE enables computation on encrypted data and offers cryptographically provable privacy. Recent work demonstrates an encryption‑friendly LLM architecture that uses LoRA fine‑tuning and Gaussian kernels to speed up fully homomorphic encryption (FHE) operations, allowing encrypted inference with performance comparable to plaintext models [gentry2009fully].
  • Trusted Execution Environments (TEEs): TEEs provide hardware‑enforced isolation that protects code and data [costan2016intel]. They partition processors into secure and normal worlds; implementations include Intel SGX, AMD SEV and Arm TrustZone. In secure LLM inference, TEEs ensure that models run only inside authenticated enclaves, with attestation verifying model integrity.
@@
4.2.5 Governance
Governance integrates ethical, legal, and societal considerations into the design and operation of AI systems. Below, we summarize how major governance frameworks inform the TRiSM principles.
  • EU AI Act — Human Oversight (Article 14): High-risk AI systems must be designed so that natural persons can effectively oversee them and prevent or minimize risks to health, safety, or fundamental rights [EU_AI_Act2025]. Oversight mechanisms should enable users to understand capabilities and limitations, detect anomalies, interpret outputs, decide not to use the system, or stop operation (e.g., via a “stop” mechanism). In certain high-risk identification contexts, decisions may require independent verification by multiple competent individuals.
  • OECD AI Principles: These intergovernmental principles (updated in 2024) promote human-centric, trustworthy AI [oecd_ai_catalogue]. They emphasize transparency and explainability: AI actors should provide meaningful information about data sources, factors, and logic so affected individuals can understand and, where appropriate, challenge outputs. Robustness, security, and safety require systems to function appropriately under normal and adverse conditions, including mechanisms to override or decommission systems to prevent harm. Accountability requires that AI actors ensure traceability of datasets and decisions and apply systematic risk management throughout the lifecycle.
  • NIST AI Risk Management Framework (AI RMF): The AI RMF and its Generative AI profile define four core functions: Govern, Map, Measure, and Manage [ai2023artificial]. For generative AI, NIST recommends documenting risks and potential impacts, applying interpretability techniques (e.g., gradient-based attributions, occlusion tests, and counterfactual prompting) to improve transparency, and establishing policies for oversight across the AI lifecycle.
  • Auditability and role-based access: AMAS should maintain auditable logs of agent actions, decisions, and data flows, and assign clear roles (e.g., reasoning, verification, summarization) so responsibilities remain transparent [khan2024role]. Role-based access control (RBAC) helps enforce accountability and prevents unauthorized agents from invoking sensitive functions.
  • Global harmonisation: Many jurisdictions (e.g., GDPR, CCPA) require data-subject rights, consent, and privacy by design. Frameworks such as the OECD Principles and NIST AI RMF also promote international cooperation and interoperability.
These governance frameworks (Figure 12) highlight the need for clear accountability structures, human oversight, robust privacy controls, and documented processes in AMAS. TRiSM provides a practical scaffolding to operationalize these requirements by aligning lifecycle management, security, privacy, and governance to support trustworthy agentic AI.
@@ 4.2.5 Governance
Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art. 13–14) [eu_ai_act] and the NIST AI RMF trustworthiness characteristic “Explainability & Interpretability” [ai2023artificial]; also supported by NISTIR 8312 (Four Principles of Explainable AI) [phillips2021four]. Promotes well-justified (rationale) and human-interpretable outputs with traceable logs [tomsett2018interpretable]. ModelOps: Anchored in the NIST AI RMF “Govern/Manage” functions [ai2023artificial] and ISO/IEC 42001:2023 (AI management system) [thiers2024emerging]; secure SDLC controls via NIST SP 800-218/218A [^9] NIST, 2024. Emphasizes lifecycle gates, model versioning, rollback policies, and CI/CD with safety gates. Risk Management & Impact Assessment: EU AI Act (Art. 9, 72–73) [eu_ai_act], ISO/IEC 23894 (AI risk management) [^10] ISO/IEC 23894, and ISO/IEC 42005:2025 (AI impact assessment) [ISO_IEC_42005_2025]. Iterative hazard analysis, impact registers, incident response, and disclosure playbooks. Data Governance & Quality: EU AI Act (Art. 10) [eu_ai_act] mapped to the ISO/IEC 5259 series (data-quality model, measures, processes, governance) [^11] ISO/IEC 5259 series. Data minimization, lineage, coverage/bias checks, and sampling documentation. Robustness & Safety: EU AI Act (Art. 15) [eu_ai_act] and ISO/IEC 24029-1 (robustness assessment) [iso24029-1]. Adversarial/perturbation testing, fault tolerance, fallback/kill-switch triggers, and rollback criteria. Application Security: Guided by OWASP LLM Top-10 [owasp2024llmtop10] and OECD AI Principles on robustness [oecd_ai_catalogue]. Access controls, prompt/input mediation, sandboxing, egress controls, and auditable trails. Model Privacy: Aligned with GDPR (Art. 25) [gdpr25], HIPAA §164 [hipaa164], and OECD AI Principles [oecd_ai_catalogue]. Encryption in transit/at rest, DP budgeting, data minimization, and access-audit pass thresholds. Documentation & Traceability: EU AI Act (Arts. 11–12; Annex IV) [eu_ai_act]. Complete technical documentation, event logs, model/data cards, decision records, and regulator-ready evidence packages. Post-Market Monitoring: EU AI Act (Arts. 72–73) [eu_ai_act]. Telemetry for drift/misuse, continuous evaluations, feedback loops, controlled recalls, and stakeholder notifications. Governance: Grounded in GDPR, CCPA [oag_ccpa_2024], and OECD AI Principles [oecd_ai_catalogue]. Institutional oversight, policy compliance, role-based accountability, regulatory traceability, and human-in-the-loop for high-risk/multi-agent deployments. Fairness & Human Rights: OECD AI Principles [oecd_ai_catalogue] and ISO/IEC TR 24027 (bias) [^12] ISO/IEC TR 24027. Lifecycle bias controls, affected-party impact reviews, and mitigation sign-offs.
Key Insights
This section operationalized the TRiSM framework for AMAS across five pillars: Explainability, ModelOps, Application Security, Model Privacy, and Governance. Multi-agent explainability requires novel techniques including Layered CoT decomposition, decision provenance graphs, and role-based interpretability frameworks to trace distributed decision processes. ModelOps extends lifecycle management through hierarchical monitoring, prompt versioning, and CI/CD safety gates with pre-deployment simulation. Security strategies emphasize defense-in-depth via design patterns (Plan-Then-Execute), least-privilege access control, and cross-agent verification. Privacy preservation leverages differential privacy, secure multi-party computation, homomorphic encryption, and trusted execution environments to protect sensitive data during inter-agent communication. Governance integration with regulatory frameworks (EU AI Act, NIST AI RMF, ISO/IEC 42001) ensures that TRiSM implementation satisfies transparency, accountability, and human oversight requirements for high-stakes deployments.
@@
5 TRiSM Evaluation of Agentic AI Systems
Accuracy alone is insufficient for Agentic AI, where failures often arise from coordination breakdowns, unsafe tool use, or misaligned autonomy rather than incorrect predictions. We organize evaluation into five dimensions: trustworthiness, explainability, user-centered performance, coordination, and composite reporting, each capturing a distinct facet of real-world performance and impact. We summarize these dimensions in Table 6.
Trustworthiness